OpenStudy (anonymous):

Two towns that are 31.5 km apart are 3 cm apart on a map. What is the scale of the map

OpenStudy (anonymous):

Scale is the ratio of distance on the map to distance on the ground. Put both measurements in the same units, then find the ratio.

OpenStudy (anonymous):

For example, if 2 cm = 20 km, I would change km to cm, so: \(20 \text{ km} \times \frac{10^5 \text{ cm}}{1 \text{ km}} = 2 \times 10^6 \text{ cm}\) \(2 \text{ cm (map)} : 2000000 \text{ cm (ground)}\) Simplify to get: 1:1000000